OpenStudy (anonymous):

-1/7 divided by 4rt3/7. rt being square root symbol. How do I divide expressions with radicals by a fraction? Help, please. :)

Parth (parthkohli):

\[{-1 \over 7} \div {4\sqrt3 \over 7 }\]Now, do remember that,\[{a \over b} \div {c \over d} = {a \div c \over b \div d}\]

Parth (parthkohli):

\[{-1 \over 7} \div {4\sqrt3 \over 7} \implies {-1 \div 4\sqrt{3} \over 7 \div 7} \implies {-1 \div 4\sqrt3}\]Can you continue?

OpenStudy (anonymous):

Would you use a reciprocal or conjugate at this point?

OpenStudy (anonymous):

Ohh, actually I would just multiply numerator and denominator by \[\sqrt{3}\] to get -\[-\frac{ \sqrt{3} }{ 12 }\]

OpenStudy (anonymous):

Thanks so much. :)